# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ 5 Concluding remarks
In this paper, we propose XVAE, a new variational autoencoder, which integrates a novel max-id model for spatial extremes that exhibits flexible extremal dependence properties. It greatly advances the ability to model extremes in high-dimensional spatial problems and expands the frontier on computation and modeling of complex extremal processes. The encoder and decoder construction and the trained distributions of the latent variables allow for parameter estimation and uncertainty quantification within a variational Bayesian framework. We also provide general guidance on evaluating emulator performance when applied to spatial data with dependent extremes.
@@
We note that our emulator extends beyond emulating large datasets for UQ. As highlighted in the introduction, the XVAE can serve as a surrogate model for mechanic-based computer models. It can also be applied to areas other than climate-related problems. For example, turbulent buoyant plume can be simulated from a system of compressible Euler conservative equations in flux formulation, but the computational cost is prohibitively expensive with increasing Reynolds number (Bhimireddy and Bhaganagar, 2021). Our XVAE can provide a promising avenue for efficiently emulating the chaotic and irregular turbulence observations in high resolutions.
@@
One possible model improvement is that the latent exponentially-tilted PS variables are independent over space and time, which is unrealistic for physical processes that exhibit diffusive dynamics at short-time scales. In future work, we are planning to include a time component with data-driven dynamic learning based on a stochastic dynamic spatio-temporal model. Hence, the latent variables in the encoded space will evolve smoothly over time while retaining heavy tails and thus simultaneously ensuring local extremal dependence. Also, it is possible to improve the XVAE by allowing spatially-varying radii $r_{k}$, $k=1,\ldots,K$, and estimate them by optimizing the ELBO together with other parameters.
@@
Another promising direction for future work is to implement a conditional VAE (CVAE; Sohn et al., 2015) with a similar underlying max-id model; in such a model, we can allow the parameters of both the encoder and decoder to change conditioning on different climate scenarios (e.g., radiative forcings, seasons, soil conditions, etc.). This will allow us to simulate new data under different conditions. We will need to ensure that the CVAE emulates $\boldsymbol{x}_{t}$ differently according to different input states (e.g., tuning parameters and/or forcing variables). In doing so, we will allow changes to the parameters for both the encoder and decoder conditioning on different scenarios (e.g., different climate states).
Appendix A Comparison to existing spatial extremes models
Max-stable (Davison et al., 2012, 2019) or generalized Pareto processes (Ferreira and de Haan, 2014; Thibaud and Opitz, 2015) have the property that $\chi_{ij}$ is always positive (Huser and Wadsworth, 2022). Conversely, Gaussian processes (or multivariate Gaussian distributions) have the property that $\chi_{ij}(u)$ always converges to 00 as $u\rightarrow 1$, unless $X_{i}$ and $X_{j}$ are perfectly dependent.
By contrast, tail dependence in observed environmental processes often seems to decay as events get more extreme and rare events often tend to be more spatially localized as the intensity increases (Huser and Wadsworth, 2019). This was observed in numerous studies, including Dutch wind gust maxima (Huser et al., 2021), threshold exceedances of the daily Fosberg fire index (Zhang et al., 2022), and winter maximum precipitation data over the Midwest of the U.S. (Zhang et al., 2023), just to name a few examples. The stability property of max-stable and generalized Pareto models is thus often a physically inappropriate restriction in the joint tail. However, a weakening $\chi_{ij}(u)$ as $u$ increases does not necessarily lead to AI. As we extrapolate into the joint tail beyond the observed data, mis-classifying the tail dependence regime inevitably leads to inaccurate risk assessments. Therefore, we seek models that exhibit much more flexible tail characteristics and do not assume an extremal a dependence class a priori.
Bopp et al. (2021) proposed a max-id model that exhibits AI ($\gamma_{k}\equiv\gamma>0$) or AD ($\gamma_{k}\equiv 0$) for all pairs of locations, depending on the tail dependence strength in the data. This is an improvement over Reich and Shaby (2012) in which the tilting parameters are set to $\gamma_{k}\equiv 0$, $k=1,\ldots,K$. In both cases, they used the radial basis functions which are not compactly supported and the tilting parameters are fixed as a single constant at all knots. As a result, these processes only exhibit one dependence class for all pairs of locations and induce a form of long-range dependence; that is, they cannot capture a change of asymptotic dependence class as a function of distance.
In addition, both Reich and Shaby (2012) and Bopp et al. (2021) set $\alpha_{0}=\alpha$, and their noise process $\{\epsilon(\boldsymbol{s})\}$ has independent $(1/\alpha)$-Fréchet marginals, i.e., Fréchet$(0,1,1/\alpha)$. While $\alpha$ also determines the dependence properties of the process through the exponentially-tilted PS variables in $\{Y(\boldsymbol{s})\}$, the Fréchet$(0,1,1/\alpha)$ variables usually end up being too noisy compared to the $\{Y(\boldsymbol{s})\}$ process.
In our modification, we decouple the noise variance and the tail heaviness of latent variables. Specifically, the Fréchet$(0,\tau,1/\alpha_{0})$ distribution concentrates around $\tau$ when the shape $11/\alpha_{0}>$ 1 and scale $\tau>1$, so that $\epsilon(\boldsymbol{s})$ truly acts as a scaling factor that accounts for measurement errors for each time $t$. Moreover, the spatially-varying tilting parameters introduce tail dependence for two close-by locations, when they are covered by the same basis function. The local dependence strength is proportional to the tail-heaviness of the latent variable at the closest knot. As we show in Section 1.3, there is local AD if $\gamma_{k}=0$, and there is local AI if $\gamma_{k}>0$, as expected. In addition, the compactness of the basis function support automatically introduces long-range exact independence (thus, also asymptotic independence) for two far-apart stations that are impacted by disjoint sets of basis functions.Overall, the dependence structure of our model is thus both non-stationary and highly flexible.
More importantly, the combination of the nonstationary max-id model with the VAE technique allows us to fit complicated spatial extremes process in exceptionally high dimensions. Beyond certain specific isolated exceptions, it is currently not possible to fit existing max-stable, inverted-max-stable, and other spatial extremes models using a full likelihood or Bayesian approach on a dataset of more than approximately $1,000$ locations. Recent successes include Huser et al. (2023), who used the Vecchia approximation to make efficient inference on a $1000$-dimensional max-stable process, Simpson et al. (2023) who used INLA to fit the spatial conditional extremes model in high dimensions, and Sainsbury-Dale et al. (2024) and Richards et al. (2023), who performed likelihood-free inference with neural Bayes estimators to fit various spatial extremes models on big data sets. However, Sainsbury-Dale et al. (2024) illustrated their method with a stationary process observed at about $700$ locations, while Richards et al. (2023) fitted local stationary models to data sets of size up to about $1,000$. By contrast, our approach considers fitting and simulating a globally non-stationary spatial extremes process, with parameters evolving over time, to a data set of unprecedented spatial dimension of more than $16,000$ locations.
Appendix B Technical details
B.1 Properties of exponentially-tilted positive-stable variables
@@
Before we proceed to prove Proposition 1.1, we first recall some useful results in Hougaard (1986) about positive-stable (PS) distributions and their exponentially-tilted variation. If $Z\sim\mathrm{expPS}(\alpha,0)$, we denote the density function by $f_{\alpha}(z)$, $z>0$. Then for $\alpha\in(0,1]$, it has Laplace transform
@@
$L(s)=\mathbb{E}{e^{-sZ}}=\exp\left[-\{(\gamma+s)^{\alpha}-\gamma^{\alpha}\}\right],\;s\geq 0,\;\gamma\geq 0$ (B.1)
@@
From Proposition 3.2 of Nolan (2020), we know that the Laplace transform of $Z\sim\text{Stable}(\alpha,1,\xi,0;1)$, $\alpha\in(0,2]$, is
@@
To facilitate the computation of the prior in Eq. (6), we follow the Monte Carlo integration steps in Section 4 of the Supplementary Material of Bopp et al. (2021) to calculate the density $h(\cdot;\alpha,\gamma)$.
B.2 Proof of Proposition 1.1
@@
To study the tail decay of the survival function, we apply Taylor’s expansion with the Peano remainder:
$(1+t)^{\alpha}=1+\alpha t+\frac{\alpha(\alpha-1)}{2}t^{2}+o(t^{2}),\;\text{as }t\rightarrow 0.$ (B.2)
@@
$\sum_{k\in\bar{\mathcal{D}}}$$\left\{\gamma_{k}+\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\omega_{kj}^{\frac{1}{\alpha}}\right\}^{\alpha}=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}\left\{1+\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\frac{\omega_{kj}^{1/\alpha}}{\gamma_{k}}\right\}^{\alpha}$
$=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}+\alpha\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\sum_{k\in\bar{\mathcal{D}}}\frac{\omega_{kj}^{1/\alpha}}{\gamma_{k}^{1-\alpha}}+\frac{\alpha(\alpha-1)}{2}\left(\frac{\tau}{x}\right)^{\frac{2}{\alpha_{0}}}\sum_{k\in\bar{\mathcal{D}}}\frac{\omega_{kj}^{2/\alpha}}{\gamma_{k}^{2-\alpha}}+o\left(x^{-\frac{2}{\alpha_{0}}}\right),$
@@
$\begin{split}\sum_{k\in\bar{\mathcal{D}}}&\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left\{\gamma_{k}+\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\omega_{kj}^{\frac{1}{\alpha}}\right\}^{\alpha}=-\left(\frac{\tau}{x}\right)^{\frac{\alpha}{\alpha_{0}}}\sum_{k\in\mathcal{D}}\omega_{kj}-\alpha\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\sum_{k\in\bar{\mathcal{D}}}\frac{\omega_{kj}^{1/\alpha}}{\gamma_{k}^{1-\alpha}}-\\ &\frac{\alpha(\alpha-1)}{2}\left(\frac{\tau}{x}\right)^{\frac{2}{\alpha_{0}}}\sum_{k\in\bar{\mathcal{D}}}\frac{\omega_{kj}^{2/\alpha}}{\gamma_{k}^{2-\alpha}}+o\left(x^{-\frac{2}{\alpha_{0}}}\right)=-c_{j}^{\prime}x^{-\frac{\alpha}{\alpha_{0}}}-c_{j}x^{-\frac{1}{\alpha_{0}}}-d_{j}x^{-\frac{2}{\alpha_{0}}}+o\left(x^{-\frac{2}{\alpha_{0}}}\right),\end{split}$ (B.3)
where the constants $c^{\prime}_{j}$, $c_{j}$ and $d_{j}$ are defined in Proposition 1.1.
Next we apply the following Taylor expansion:
$1-\exp(-t)=t-\frac{t^{2}}{2}+o(t^{2}),\;\text{as }t\rightarrow 0$ (B.4)
Combining (10) and (B.3) gives
$\bar{F}_{j}(x)=c_{j}^{\prime}x^{-\frac{\alpha}{\alpha_{0}}}+c_{j}x^{-\frac{1}{\alpha_{0}}}+d_{j}x^{-\frac{2}{\alpha_{0}}}-\frac{(c_{j}^{\prime}x^{-\frac{\alpha}{\alpha_{0}}}+c_{j}x^{-\frac{1}{\alpha_{0}}}+d_{j}x^{-\frac{2}{\alpha_{0}}})^{2}}{2}+o\left(x^{-\frac{2}{\alpha_{0}}}\right),$
from which we can expand the squared term and discard the terms with higher decaying rates than $o(x^{-2/\alpha_{0}})$ to establish (11).
Proof of Corollary 1.1.1.
By definition, $t^{-1}=\bar{F}_{j}\{q_{j}(t)\}$. When $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, (11) leads to
$\begin{split}t^{-1}=c^{\prime}_{j}q_{j}^{-\frac{\alpha}{\alpha_{0}}}&(t)\left[1+\frac{c_{j}}{c^{\prime}_{j}}q_{j}^{-\frac{1-\alpha}{\alpha_{0}}}(t)+\frac{1}{c^{\prime}_{j}}\left(d_{j}-\frac{c_{j}^{2}}{2}\right)q_{j}^{-\frac{2-\alpha}{\alpha_{0}}}(t)-\right.\\ &\left.\frac{c^{\prime}_{j}}{2}q_{j}^{-\frac{\alpha}{\alpha_{0}}}(t)-c_{j}q_{j}^{-\frac{1}{\alpha_{0}}}(t)+o\left\{q_{j}^{-\frac{2-\alpha}{\alpha_{0}}}(t)\right\}\right]\;\text{as }t\rightarrow\infty.\end{split}$ (B.5)
@@
$q_{j}(t)=c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}\{1+o(1)\}.$ (B.6)
Since $\alpha\in(0,1)$, we can also re-organize (B.5) to obtain
$\begin{split}q_{j}(t)-c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}&=q_{j}(t)\left(1-\left[1+\frac{c_{j}}{c^{\prime}_{j}}q_{j}^{-\frac{1-\alpha}{\alpha_{0}}}(t)-\frac{c^{\prime}_{j}}{2}q_{j}^{-\frac{\alpha}{\alpha_{0}}}(t)+O\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right]^{-\frac{\alpha_{0}}{\alpha}}\right)\\ &=q_{j}(t)\left[\frac{\alpha_{0}c_{j}}{\alpha c^{\prime}_{j}}q_{j}^{-\frac{1-\alpha}{\alpha_{0}}}(t)-\frac{\alpha_{0}c^{\prime}_{j}}{2\alpha}q_{j}^{-\frac{\alpha}{\alpha_{0}}}(t)+O\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right].\end{split}$ (B.7)
On the last line, we applied the Taylor expansion in (B.2) again. Then we combine (B.6) and (B.7) to get
$q_{j}(t)-c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}$ $=c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}\{1+o(1)\}\left\{\frac{\alpha_{0}c_{j}}{\alpha c^{\prime}_{j}{}^{1/\alpha}}t^{1-\frac{1}{\alpha}}-\frac{\alpha_{0}}{2\alpha}t^{-1}+O\left(t^{-\frac{1}{\alpha}}\right)\right\}$
$=c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}\left\{\frac{\alpha_{0}c_{j}}{\alpha c^{\prime}_{j}{}^{1/\alpha}}t^{1-\frac{1}{\alpha}}-\frac{\alpha_{0}}{2\alpha}t^{-1}+O\left(t^{-\frac{1}{\alpha}}\right)\right\},$
@@
$c_{j}^{\alpha_{0}}t^{\alpha_{0}}=q_{j}(t)\left[1+\left(\frac{d_{j}}{c_{j}}-\frac{c_{j}}{2}\right)q_{j}^{-\frac{1}{\alpha_{0}}}(t)+o\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right]^{-\alpha_{0}}\;\text{as }t\rightarrow\infty,$
@@
$q_{j}(t)-c_{j}^{\alpha_{0}}t^{\alpha_{0}}$ $=q_{j}(t)\left(1-\left[1+\left(\frac{d_{j}}{c_{j}}-\frac{c_{j}}{2}\right)q_{j}^{-\frac{1}{\alpha_{0}}}(t)+o\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right]^{-\alpha_{0}}\right)$
$=c_{j}^{\alpha_{0}}t^{\alpha_{0}}\{1+o(1)\}\left[\alpha_{0}\left(\frac{d_{j}}{c_{j}}-\frac{c_{j}}{2}\right)q_{j}^{-\frac{1}{\alpha_{0}}}(t)+o\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right]$
$=c_{j}^{\alpha_{0}}t^{\alpha_{0}}\left\{\alpha_{0}\left(\frac{d_{j}}{c_{j}^{2}}-\frac{1}{2}\right)t^{-1}+o(t^{-1})\right\}.$
B.3 Proof of Proposition 1.2
@@
in which we utilized the Laplace transform of the exponentially-tilted PS variables displayed in Eq. (B.1). ∎
B.4 Proof of Theorem 1.3

### v4

@@ 6 Concluding remarks
In this paper, we propose XVAE, a new variational autoencoder, which integrates a novel max-id model for spatial extremes that exhibits flexible extremal dependence properties. It greatly advances the ability to model extremes in high-dimensional spatial problems and expands the frontier on computation and modeling of complex extremal processes. The encoder and decoder construction and the trained distributions of the latent variables allow for parameter estimation and uncertainty quantification within a variational Bayesian framework. We also develop a validation framework for evaluating emulator performance when applied to spatial data with dependent extremes.
@@
We note that our emulator extends beyond emulating large datasets for UQ. As highlighted in the introduction, the XVAE can serve as a surrogate model for mechanistic-based computer models. It can also be applied to areas other than climate-related problems. For example, turbulent buoyant plume can be simulated from a system of compressible Euler conservative equations in flux formulation, but the computational cost is prohibitively expensive with increasing Reynolds number (Bhimireddy and Bhaganagar, 2021). Our XVAE can provide a promising avenue for efficiently emulating the chaotic and irregular turbulence observations at high resolutions.
@@
One major drawback of XVAE is that the latent expPS variables are independent over space and time, which is unrealistic for physical processes that exhibit dynamics at short-time scales. As a result, it cannot capture temporal dependence appropriately. In future work, we are planning to include a time component with data-driven dynamic learning based on a stochastic dynamic spatio-temporal model. Hence, the latent variables in the encoded space will evolve smoothly over time while retaining heavy tails and thus simultaneously ensuring local extremal dependence. Furthermore, it is possible to improve the XVAE by allowing the basis functions’ radii $r_{k}$, $k=1,\ldots,K$, to be spatially-varying, and estimating them by optimizing the ELBO together with the other parameters.
@@
Another promising direction for future work is to implement a conditional VAE (CVAE; Sohn et al., 2015) with a similar underlying max-id model; in such a model, we can allow the parameters of both the encoder and decoder to change conditional on different climate scenarios (e.g., radiative forcings, seasons, soil conditions, etc.). This will allow us to simulate new data under different conditions. We will need to ensure that the CVAE emulates $\boldsymbol{x}_{t}$ differently according to different input states (e.g., tuning parameters and/or forcing variables). In doing so, we will allow changes to the parameters for both the encoder and decoder conditioning on different scenarios (e.g., different climate states).
Appendix A Technical details
A.1 Properties of exponentially-tilted positive-stable variables
@@
Before we proceed to prove Proposition 3.1 from the main paper, we first recall some useful results in Hougaard (1986) about positive-stable (PS) distributions and their exponentially-tilted variation. If $Z\sim\mathrm{expPS}(\alpha,0)$, we denote the density function by $f_{\alpha}(z)$, $z>0$. Then for $\alpha\in(0,1]$, it has Laplace transform
@@
$L(s)=\mathbb{E}{e^{-sZ}}=\exp\left[-\{(\gamma+s)^{\alpha}-\gamma^{\alpha}\}\right],\;s\geq 0,\;\gamma\geq 0$ (A.1)
@@
From Proposition 3.2 of Nolan (2020), we know that the Laplace transform of the random variable $Z\sim\text{Stable}(\alpha,1,\xi,0;1)$, $\alpha\in(0,2]$, is
@@
To facilitate the computation of the prior in Eq. (9) of the main paper, we follow the Monte Carlo integration steps in Section 4 of the Supplementary Material of Bopp et al. (2021) to calculate the density $h(\cdot;\alpha,\gamma)$.
A.2 Proof of Proposition 3.1 of the main paper
@@
We now show that as $x\rightarrow\infty$, the survival function $\bar{F}_{j}(x)=1-F_{j}(x)$ satisfies
$\bar{F}_{j}(x)=c^{\prime}_{j}\left(\frac{x}{\tau}\right)^{-\frac{\alpha}{\alpha_{0}}}+c_{j}\left(\frac{x}{\tau}\right)^{-\frac{1}{\alpha_{0}}}+\left(d_{j}-\frac{c_{j}^{2}}{2}\right)\left(\frac{x}{\tau}\right)^{-\frac{2}{\alpha_{0}}}-\frac{c^{\prime}_{j}{}^{2}}{2}\left(\frac{x}{\tau}\right)^{-\frac{2\alpha}{\alpha_{0}}}-c^{\prime}_{j}c_{j}\left(\frac{x}{\tau}\right)^{-\frac{\alpha+1}{\alpha_{0}}}+o\left(x^{-\frac{2}{\alpha_{0}}}\right),$ (A.2)
where $c_{j}=\alpha\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha-1}\omega_{kj}^{1/\alpha}$, $c^{\prime}_{j}=\sum_{k\in\mathcal{D}}\omega_{kj}$, and $d_{j}=\frac{\alpha(\alpha-1)}{2}\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha-2}\omega_{kj}^{2/\alpha}$.
First, we apply Taylor’s expansion with the Peano remainder:
$(1+t)^{\alpha}=1+\alpha t+\frac{\alpha(\alpha-1)}{2}t^{2}+o(t^{2}),\;\text{as }t\rightarrow 0.$ (A.3)
@@
$\begin{split}\sum_{k\in\bar{\mathcal{D}}}&\left\{\gamma_{k}+\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\omega_{kj}^{\frac{1}{\alpha}}\right\}^{\alpha}=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}\left\{1+\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\frac{\omega_{kj}^{1/\alpha}}{\gamma_{k}}\right\}^{\alpha}\\ &=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}+\alpha\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\sum_{k\in\bar{\mathcal{D}}}\frac{\omega_{kj}^{1/\alpha}}{\gamma_{k}^{1-\alpha}}+\frac{\alpha(\alpha-1)}{2}\left(\frac{\tau}{x}\right)^{\frac{2}{\alpha_{0}}}\sum_{k\in\bar{\mathcal{D}}}\frac{\omega_{kj}^{2/\alpha}}{\gamma_{k}^{2-\alpha}}+o\left(x^{-\frac{2}{\alpha_{0}}}\right),\end{split}$
@@
$\begin{split}\sum_{k\in\bar{\mathcal{D}}}&\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left\{\gamma_{k}+\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\omega_{kj}^{\frac{1}{\alpha}}\right\}^{\alpha}=-\left(\frac{\tau}{x}\right)^{\frac{\alpha}{\alpha_{0}}}\sum_{k\in\mathcal{D}}\omega_{kj}-\alpha\left(\frac{\tau}{x}\right)^{\frac{1}{\alpha_{0}}}\sum_{k\in\bar{\mathcal{D}}}\frac{\omega_{kj}^{1/\alpha}}{\gamma_{k}^{1-\alpha}}\\ &-\frac{\alpha(\alpha-1)}{2}\left(\frac{\tau}{x}\right)^{\frac{2}{\alpha_{0}}}\sum_{k\in\bar{\mathcal{D}}}\frac{\omega_{kj}^{2/\alpha}}{\gamma_{k}^{2-\alpha}}+o\left(x^{-\frac{2}{\alpha_{0}}}\right)=-c_{j}^{\prime}\left(\frac{x}{\tau}\right)^{-\frac{\alpha}{\alpha_{0}}}-c_{j}\left(\frac{x}{\tau}\right)^{-\frac{1}{\alpha_{0}}}-d_{j}\left(\frac{x}{\tau}\right)^{-\frac{2}{\alpha_{0}}}+o\left(x^{-\frac{2}{\alpha_{0}}}\right),\end{split}$
where the constants $c^{\prime}_{j}$, $c_{j}$ and $d_{j}$ are defined in Proposition 3.1 from the main paper.
Next we apply the following Taylor expansion
$1-\exp(-t)=t-\frac{t^{2}}{2}+o(t^{2}),\;\text{as }t\rightarrow 0$ (A.4)
to get
$\bar{F}_{j}$ $(x)=c_{j}^{\prime}\left(\frac{x}{\tau}\right)^{-\frac{\alpha}{\alpha_{0}}}+c_{j}\left(\frac{x}{\tau}\right)^{-\frac{1}{\alpha_{0}}}+d_{j}\left(\frac{x}{\tau}\right)^{-\frac{2}{\alpha_{0}}}$
$-\frac{1}{2}\left\{c_{j}^{\prime}\left(\frac{x}{\tau}\right)^{-\frac{\alpha}{\alpha_{0}}}+c_{j}\left(\frac{x}{\tau}\right)^{-\frac{1}{\alpha_{0}}}+d_{j}\left(\frac{x}{\tau}\right)^{-\frac{2}{\alpha_{0}}}\right\}^{2}+o\left(x^{-\frac{2}{\alpha_{0}}}\right),$
from which we can expand the squared term and discard the terms with higher decaying rates than $o(x^{-2/\alpha_{0}})$ to establish (A.2).
Lastly, from (A.2), it is clear that as $x\rightarrow\infty$, $\bar{F}_{j}(x)\sim c_{j}(x/\tau)^{-{1/\alpha_{0}}}$ if $\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$, and $\bar{F}_{j}(x)\sim c_{j}^{\prime}(x/\tau)^{-{\alpha/\alpha_{0}}}$ if $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$.
The following result directly delineates how the quantile level changes as $u\rightarrow 1$. It will be used to derive the tail dependence structure for two arbitrary spatial locations.
Corollary A.1.1.
As $t\rightarrow\infty$, the marginal quantile function $q_{j}(t)=F_{j}^{-1}(1-1/t)$ can be approximated as follows under the assumptions of Proposition 3.1 from the main paper:
$q_{j}(t)=\begin{cases}\tau c^{\prime}_{j}{}^{\alpha_{0}/\alpha}t^{\alpha_{0}/\alpha}\left\{1+\frac{\alpha_{0}c_{j}t^{1-1/\alpha}}{\alpha c^{\prime}_{j}{}^{1/\alpha}}-\frac{\alpha_{0}t^{-1}}{2\alpha}+O\left(t^{-1/\alpha}\right)\right\},&\text{if }\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset,\\ \tau c_{j}^{\alpha_{0}}t^{\alpha_{0}}\{1+\alpha_{0}(\frac{d_{j}}{c_{j}^{2}}-\frac{1}{2})t^{-1}+o(t^{-1})\},&\text{if }\mathcal{C}_{j}\cap\mathcal{D}=\emptyset.\end{cases}$
Proof.
By definition, $t^{-1}=\bar{F}_{j}\{q_{j}(t)\}$. When $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, (A.2) leads to
$\begin{split}t^{-1}=c^{\prime}_{j}\tau^{\frac{\alpha}{\alpha_{0}}}q_{j}^{-\frac{\alpha}{\alpha_{0}}}&(t)\left[1+\frac{c_{j}\tau^{\frac{1-\alpha}{\alpha_{0}}}}{c^{\prime}_{j}}q_{j}^{-\frac{1-\alpha}{\alpha_{0}}}(t)+\frac{c_{j}\tau^{\frac{2-\alpha}{\alpha_{0}}}}{c^{\prime}_{j}}\left(d_{j}-\frac{c_{j}^{2}}{2}\right)q_{j}^{-\frac{2-\alpha}{\alpha_{0}}}(t)-\right.\\ &\left.\frac{c^{\prime}_{j}\tau^{\frac{\alpha}{\alpha_{0}}}}{2}q_{j}^{-\frac{\alpha}{\alpha_{0}}}(t)-c_{j}\tau^{\frac{1}{\alpha_{0}}}q_{j}^{-\frac{1}{\alpha_{0}}}(t)+o\left\{q_{j}^{-\frac{2-\alpha}{\alpha_{0}}}(t)\right\}\right]\;\text{as }t\rightarrow\infty.\end{split}$ (A.5)
@@
$q_{j}(t)=\tau c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}\{1+o(1)\}.$ (A.6)
Since $\alpha\in(0,1)$, we can also re-organize (A.5) to obtain
$\begin{split}q_{j}(t)-\tau c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}&=q_{j}(t)\left(1-\left[1+\frac{c_{j}\tau^{\frac{1-\alpha}{\alpha_{0}}}}{c^{\prime}_{j}}q_{j}^{-\frac{1-\alpha}{\alpha_{0}}}(t)-\frac{c^{\prime}_{j}\tau^{\frac{\alpha}{\alpha_{0}}}}{2}q_{j}^{-\frac{\alpha}{\alpha_{0}}}(t)+O\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right]^{-\frac{\alpha_{0}}{\alpha}}\right)\\ &=q_{j}(t)\left[\frac{\alpha_{0}c_{j}\tau^{\frac{1-\alpha}{\alpha_{0}}}}{\alpha c^{\prime}_{j}}q_{j}^{-\frac{1-\alpha}{\alpha_{0}}}(t)-\frac{\alpha_{0}c^{\prime}_{j}\tau^{\frac{\alpha}{\alpha_{0}}}}{2\alpha}q_{j}^{-\frac{\alpha}{\alpha_{0}}}(t)+O\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right].\end{split}$ (A.7)
On the last line, we applied the Taylor expansion in (A.3) again. Then we combine (A.6) and (A.7) to get
$q_{j}(t)-\tau c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}$ $=\tau c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}\{1+o(1)\}\left\{\frac{\alpha_{0}c_{j}}{\alpha c^{\prime}_{j}{}^{1/\alpha}}t^{1-\frac{1}{\alpha}}-\frac{\alpha_{0}}{2\alpha}t^{-1}+O\left(t^{-\frac{1}{\alpha}}\right)\right\}$
$=\tau c^{\prime}_{j}{}^{\frac{\alpha_{0}}{\alpha}}t^{\frac{\alpha_{0}}{\alpha}}\left\{\frac{\alpha_{0}c_{j}}{\alpha c^{\prime}_{j}{}^{1/\alpha}}t^{1-\frac{1}{\alpha}}-\frac{\alpha_{0}}{2\alpha}t^{-1}+O\left(t^{-\frac{1}{\alpha}}\right)\right\},$
@@
$\tau c_{j}^{\alpha_{0}}t^{\alpha_{0}}=q_{j}(t)\left[1+\left(\frac{d_{j}}{c_{j}}-\frac{c_{j}}{2}\right)\tau^{\frac{1}{\alpha_{0}}}q_{j}^{-\frac{1}{\alpha_{0}}}(t)+o\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right]^{-\alpha_{0}}\;\text{as }t\rightarrow\infty,$
@@
$q_{j}(t)-\tau c_{j}^{\alpha_{0}}t^{\alpha_{0}}$ $=q_{j}(t)\left(1-\left[1+\left(\frac{d_{j}}{c_{j}}-\frac{c_{j}}{2}\right)\tau^{\frac{1}{\alpha_{0}}}q_{j}^{-\frac{1}{\alpha_{0}}}(t)+o\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right]^{-\alpha_{0}}\right)$
$=\tau c_{j}^{\alpha_{0}}t^{\alpha_{0}}\{1+o(1)\}\left[\alpha_{0}\left(\frac{d_{j}}{c_{j}}-\frac{c_{j}}{2}\right)\tau^{\frac{1}{\alpha_{0}}}q_{j}^{-\frac{1}{\alpha_{0}}}(t)+o\left\{q_{j}^{-\frac{1}{\alpha_{0}}}(t)\right\}\right]$
$=\tau c_{j}^{\alpha_{0}}t^{\alpha_{0}}\left\{\alpha_{0}\left(\frac{d_{j}}{c_{j}^{2}}-\frac{1}{2}\right)t^{-1}+o(t^{-1})\right\}.$
A.3 Proof of Proposition 3.2 of the main paper
@@
in which we utilized the Laplace transform of the exponentially-tilted PS variables displayed in Eq. (A.1). ∎
A.4 Proof of Theorem 3.3 of the main paper